OpenStudy (anonymous):

Convert each equation from slope-intercept form to standard form... y= -1/2x - 3

OpenStudy (anonymous):

lol....this is a standard form.... y = mx + c

OpenStudy (anonymous):

You wouldn't add -1/2x to both sides ?

OpenStudy (anonymous):

no...need...

OpenStudy (anonymous):

Okay . thanks again , lol .

OpenStudy (anonymous):

Standard form is defined as Ax+Bx=C where A is not negative and A and B are not fractional. So this should be: 1/2x+y=-3 x+2y=-6